1. Find T6 (x) in explicit polynomial form and plot it on [-1,1],[-1.2,1.2] and I-2,2].
Answer #1

70 2 3 3 3 거 56. -2,2now plot on [-1,1]

plot 32x6-48 x4 +18x2-1 x=-1 to 1 Plot: 1.0 0.5 -0.5 0.5 .0 0.5 1.

Plot on [-1.2,1.2]

lot 32x6-48 x4 +18x2-1 х=-1.2 to 1.2 Plot: -1 -0.5 0.5 .0and

Plot on [-2,2]

plot 32x6-48 x4 + 18x2-1 | x=-2to2 Plot: 800 600 400 200 -2